OpenStudy (anonymous):

Evaluate the function f(x)=-3x+9 at the indicated values. Find f(2) Find f(-2)

OpenStudy (anonymous):

1st equation: x = 3 (-3*2=-6+9=3 2nd equation: x = 15 (-3*-2=6+9=15)

OpenStudy (anonymous):

F(2)= -3(2)+9 F(-2)= -3(-2)+9 Solve. The answers are 3 and 15